OpenStudy (anonymous):

Help! Solve the following formula for the specified variable : L=a+(b-1)w for b

OpenStudy (anonymous):

hi. you can start this problem by subtraction of a from both sides.

OpenStudy (anonymous):

I tried that but then I got stuck.

OpenStudy (anonymous):

ok. what was the result of that step? something like L-a=(b-1)w?

OpenStudy (anonymous):

yes that's what I got.

OpenStudy (anonymous):

and then do I divide both sides by w? I'm not sure

OpenStudy (anonymous):

perfect. you are on the right track.

OpenStudy (anonymous):

yes. divide both sides by w. what do you get?

OpenStudy (anonymous):

So I got L-a/w = (b-1). So do I add 1 now?

OpenStudy (anonymous):

It doesn't look right.

OpenStudy (anonymous):

just to be clear, it should be (L-a)/w=b-1. the parentheses indicate that w divides L-a and just a. ok? It is a matter of good form.

OpenStudy (anonymous):

then you add 1....and voila! you have done the problem.

OpenStudy (anonymous):

so what do you get for the final answer?

OpenStudy (anonymous):

So I got L-a/w +1 as the final answer. Is that right?

OpenStudy (anonymous):

did you see what i wrote above about the parentheses? The answer is (L-a)/w+1=b. Overall, nice work PhiBee, or is it Jay. lol

OpenStudy (anonymous):

Thank you! And it's PhiBee :)

OpenStudy (anonymous):

well, it is a pleasure PhiBee. You did well. :) Is there anything else with which I can be of service?

OpenStudy (anonymous):

Well I do have one more similar question?

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

The question asked to solve for b1 in the equation N=b1-b2/g

OpenStudy (anonymous):

ok. I know that books can make a straight fraction line. Does the fraction line extend beneath both b1 and b2. \[\frac{ b1-b2 }{ g }\]...like this?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so what do you think your first step might be?

OpenStudy (anonymous):

Can I multiply by g to get rid of it?

OpenStudy (anonymous):

absolutely. if you multiply by g you are no longer dealing with a fraction. makes it sooo much easier.

OpenStudy (anonymous):

So I have to multiply g on both sides?

OpenStudy (anonymous):

yes... the rules of algebra state that what you do to one side in solving an equation, you must do to the other side as well. so what was the result of that operation?

OpenStudy (anonymous):

I got N x g = b1- b2

OpenStudy (anonymous):

perfect! you have this thing in hand. Now, the last step is....?

OpenStudy (anonymous):

I add b2?

OpenStudy (anonymous):

right again... and your answer is...?

OpenStudy (anonymous):

Ng+b2 = b1 ?

OpenStudy (anonymous):

you got it... nicely done PhiBee :)

OpenStudy (anonymous):

Yay! Thank you!

OpenStudy (anonymous):

you are welcome. any other problems giving you a little doubt?

OpenStudy (anonymous):

There are two more I might need help with if that's ok?

OpenStudy (anonymous):

:) sure. fire away

OpenStudy (anonymous):

Ok. The first question asks to solve the following formula for the specified variable: A=1/2 pi n^2 f for f. Sorry i don't have the pi sign.

OpenStudy (anonymous):

so what do you think your first move is here?

OpenStudy (anonymous):

Can I multiply by 2?

OpenStudy (anonymous):

yes... now let me be clear. the problem is 1/2 times pi times n^2 times f. None of the variables appear in a denominator. right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then multiply away (by 2 of course). :))

OpenStudy (anonymous):

So I get 2A = pi n^2 f ?

OpenStudy (anonymous):

right PhiBee. now the rest is a product of pi and n^2 and f. you can get to the answer in two steps or one.

OpenStudy (anonymous):

do you see what I mean?

OpenStudy (anonymous):

So can i divide by pi?

OpenStudy (anonymous):

yes. that'll get you there in two steps.

OpenStudy (anonymous):

So I would get 2A/pi = n^2 f ?

OpenStudy (anonymous):

looking good so far...

OpenStudy (anonymous):

and the last step (the envelope please, if you watch the academy awards) is....

OpenStudy (anonymous):

Ok I'm stuck because I want to get rid of n^2. How do I do that?

OpenStudy (anonymous):

how would you get rid of it if it was just n?

OpenStudy (anonymous):

I can't square root it can I ?

OpenStudy (anonymous):

no, you dont need to square root here.

OpenStudy (anonymous):

how would you solve this if it werent n^2, just n?

OpenStudy (anonymous):

I would divide both sides by n.

OpenStudy (anonymous):

right. so now how would you isolate f from n^2?

OpenStudy (anonymous):

give me a guess. your best guess

OpenStudy (anonymous):

I divide by f?

OpenStudy (anonymous):

no wait that was wrong....

OpenStudy (anonymous):

PhiBee, you are solving for f.

OpenStudy (anonymous):

:) good that you caught that

OpenStudy (anonymous):

you know that you can divide by any power of a variable, safely and securely. lol

OpenStudy (anonymous):

This power is really aggravating me. How do I get rid of it?!

OpenStudy (anonymous):

you dont get rid of the power. it is ok to divide by n^2.

OpenStudy (anonymous):

look at what i wrote above. you can divide by any power of a variable.

OpenStudy (anonymous):

Oh wait it's ok? That was my first guess the whole time!!! i just thought it wasn't possible!

OpenStudy (anonymous):

you are solving for f, not for n. no square root need be taken. Remember which variable you are solving for. :) now what did you get for an answer?

OpenStudy (anonymous):

So then i would FINALLY get 2A/ pi n^2. PLEASE TELL ME I GOT IT CORRECT! I will go and cry if I don't. :(

OpenStudy (anonymous):

you got it!! you are fine.

OpenStudy (anonymous):

Ok. I think i can breathe now lol

OpenStudy (anonymous):

While i am helping you, I wont let you fail...:) so breathe easy.

OpenStudy (anonymous):

the next problem is....

OpenStudy (anonymous):

I am eternally grateful!

OpenStudy (anonymous):

Oh ok now time to stop breathing again....the problem said to solve: S=p/3 (c +L)

OpenStudy (anonymous):

PhiBee, keep breathing. also, did it tell you what variable to solve for?

OpenStudy (anonymous):

Oh! Yeah it said solve for L. My lack of breathing made me forget haha

OpenStudy (anonymous):

lol. Let me ask you something. During tests, do you sometimes blank an answer, even tho you know what to do?

OpenStudy (anonymous):

Are you kidding me? All the time! my mind gets all fuzzy and I feel like I'm going to faint! :(